OpenStudy (anonymous):

Solve the system of equations. Enter your answer as an ordered pair. 5x - y = 27 5x + 4y = 42

OpenStudy (anonymous):

(6,3)

OpenStudy (anonymous):

6x + 4y = 32 -6x + 4y = 8

OpenStudy (anonymous):

-5y+y=-27 5y+4y=42 >>>5y=15 > y=3

OpenStudy (anonymous):

and x= 6

hero (hero):

lol...apex sucks....actually, I saw apex..I think it's pretty cool actually

hero (hero):

Solve for y, set both equations equal to each other....solve for x, then find y

OpenStudy (anonymous):

y=3 and x=6